which term of the g.p 1/2,3/4,9/8,... is 3^9/2^10
Answer by richwmiller(17219)

Tn = T1 * r^(n - 1)
r=3/2
3^9/2^10=1/2*3/2^(n-1)
3^9*2/2^10=3/2^(n-1)
3^9/2^9=3/2^(n-1)
(3/2)^9=3/2^(n-1)
9=n-1
10=n